A bowl contains five green olives and eight black ones. Wilma picks one olive at random and eats it. She then picks another olive. What is the probability that she gets one green and one black olive?
Answer by stanbon(75887)

A bowl contains five green olives and eight black ones. Wilma picks one olive at random and eats it. She then picks another olive. What is the probability that she gets one green and one black olive?
-----------------------------
P(2 black) = 8C2/13C2 = 0.3590
---
P(2 green) = 5C2/13C2 = 0.1282
-----------------------------------
---
P(one green and one black) = 1- [0.3590+0.1282]
----
= 1 - 0.4872
----
= 0.5128
=============================
Cheers,
Stan H.
===================